You are to create an Excel spreadsheet that calculates the issue price of a bond. Your spreadsheet should prompt the user for all relevant inputs, then calculate and clearly display the issue price. Email your completed spreadsheet to me. If you do this project with a colleague, please include both names on the spreadsheet.
Answer #1
Please provide Inputs in the following fields as per instructions
Bond Face Value "a" a
Coupon Rate offered by Bond "x" x
Frequency of interest payment per year n
Assumed Market Rate "r" r
No. of years left for redemption "n" N
Solution Part
Present Value
Present value of Redemption value of Bond after "n" years (1) =a/(1+(r/n))^(N*n)
Present Value of Interest Received in each year till redemption (2) =PV((r/n),(N*n),(a*x))
Issue Price of the Bond =(1)+(2)